OpenStudy (anonymous):

Find the slope of the following line. y=-1 A. The slope of the line is B. The slope of the line is undefined

OpenStudy (anonymous):

Really?

OpenStudy (anonymous):



OpenStudy (nirmalnema):

ax+by+c=0 slope= -a/b y=-1 y+1=0 0/1 =0 slope is zero

OpenStudy (anonymous):

And any line parallel to x-axis is having a slope of \(0\).. :)

OpenStudy (anonymous):

As your given equation is more in the form of Slope-Intercept, so compare your given equation with this \(y = mx + c\) Like this :